What volume of carbon(IV)oxide is produced at s.t.p when 7.5g of calcium trioxocarbonate(IV) reacts with excess hydrochloric acid
1
Expert's answer
2020-05-07T14:15:27-0400

CaCO3 + 2HCl = CaCl2 + H2O + CO2

n(CaCO3) = n(CO2) = "{\frac {7.5}{40+12+48}}={\frac {7.5}{100}}=0.075" mol.

At s.t.p. we can use The value of molar volume which is equal to 22.4 l/mol.

V(CO2) = 0.075*22.4 = 1.68 l